(b) What is the value of x after each of these statements is encountered in a computer program,  if x =3 before the statement is reached?

(i)  if x +2=5 then x = 3*x +5

(ii) if (x +1=4) OR (2x +2=3) then x = x +1

(iii) if (2x +3=5) AND (3x +4=13) then x = x^2 +x+7

     (iv) if (x +1=4) OR (x +2=5) then x = x +1 



1
Expert's answer
2021-04-25T17:15:39-0400

(i)

3+2=5 is true, so x=3*3+5=14

Answer: x=14


(ii)

3+1=4 is true, so (3+1=4 or 6+2-3) is true, so x=3+1=4

Answer: x=4


(iii)

2*3+3=5 if false, so (2*3+3=5 and 3*3+4=13) is false, so x=3 as at the beginning of the program

Answer: x=3


(iv)

3+1=4 is true, 3+2=5 is true, so (3+1=4 or 3+2=5) is true, so x=3+1=4

Answer: x=4